John Electronics Inc. started its business in 2019 and sold electronics worth $185,000 in 2019. The company provided a 2- year limited warranty for all sales. As per company estimate warranty costs would be 1% of sales in the first year and 3% of sales in the second year. By end of 2019 John Electronics Inc had already spent $1,510 on warranty repairs. Prepare all journal entries related to the warranty for 2019. Also d on the Balance
Answer #1

Answer:

Date Accounts Title & Explanation Debit ($) Credit ($)
31,Dec,2019 Warranty Expense A/c Dr 1,510
To Cash A/c 1,510
(Given 1% of Sales = 1,510)
Warranty Expense A/c Dr 6,040
To Estimated Warranty Liability A/c 6,040
(1% + 3% of Sales)

Note:

Estimated Warranty Expense

for 1st Year given 1% of Sales = 1,510

then total sales would be 1510 * 100 = 151,000

then for 2nd Year 3 % of Sales = 151,000 * 3%

= 4,530

Therefore, Total Estimated Warranty Liability = 1,510 + 4,530

= 6,040